OpenStudy (anonymous):

A gym in a high school is getting a new floor over the summer. The gym is square and has an area of 675 squared. What is the approximate length of the gym rounded to the nearest tenth? And If the gym tiles are $2.75 per square foot, how much will it cost to cover the entire gym floor?

OpenStudy (anonymous):

do you mean 675 feet squared?

OpenStudy (anonymous):

Yeah! Haha

OpenStudy (anonymous):

great! first part: Area = side length^2 for a square, and you want to know what s equals. A = s^2 s = sqrt A s = sqrt 675 *calculator magic* s = 25.98 = 26.0 (rounded to the nearest tenth) second part: the entire gym floor is 675 square feet. tiles cost $2.75 per square foot. 675 * 2.75 = $1856.25

OpenStudy (anonymous):

Thanks for the second part! But I'm still a little bit confused with the first half.

OpenStudy (anonymous):

What is sqrt?

OpenStudy (anonymous):

square root it's an abbreviation.. sorry for being confusing (:

OpenStudy (anonymous):

Its ok! Thanks!

OpenStudy (anonymous):

sqrt A? Huh?

OpenStudy (anonymous):

take the square root of the Area square root (675) = 25.988

OpenStudy (anonymous):

Ok! Thank you so much!